A glass in a fixed container is at a pressure of 4 atm and a temperature of 27 0C. What will be the pressure if it is heated to a temperature of 177 0 C? 



1
Expert's answer
2022-04-06T15:34:05-0400

p1/T1 = p2/T2

4 atm / (27 + 273) K = p2 / (177 + 273) K

p2 = 6 atm